# FormGym: Doing Paperwork with Agents

### v1

@@ Abstract
Completing paperwork is a challenging and time-consuming problem. Form filling is especially challenging in the pure-image domain without access to OCR, typeset PDF text, or a DOM. For computer agents, it requires multiple abilities, including multi-modal understanding, information retrieval, and tool-use. We present a novel form-filling benchmark consisting of 432 fields spread across 55 documents and 3 tasks, requiring knowledge of 236 features per user. We find that baseline VLAs achieve less than 1% accuracy in most cases, primarily due to poor localization ability. GUI agents also struggle, scoring between 10.6-68.0% despite high cost and latency. Therefore, we also contribute FieldFinder, a tool to assist LLMs in identifying where to place text on a form. With FieldFinder, all models achieve equal or better performance in all six study conditions, with a maximum increase from 2% to 56%.
@@
Matthew Toles1, Rattandeep Singh1, Isaac Song2,3, Zhou Yu1,3, 1Columbia University, 2Georgia Institute of Technology, 3Arklex.ai, Correspondence: mt3639@columbia.edu
@@ 1 Introduction
Filling out paperwork is a pervasive and tedious task. Although some paper forms have been replaced by fillable rich-text PDFs, many are only available as pure images either in their original format or as scanned physical documents. These forms represent the most challenging task because agents can only interact with the document as an image rather than the information-rich DOM or PDF typeset text and vector graphics. This task builds on prior work on document understanding, OCR, localization, and agentic workflows to evaluate end-to-end image manipulation accuracy.
In this work, we propose a new benchmark for evaluating the ability of general-purpose vision-language agents (VLAs) to perform end-to-end form completion. Our evaluation focuses on realistic use cases where an agent must interpret a document and populate fields based on a user profile. Relevant user information is provided as raw text, a SQL database, or other completed forms containing partially overlapping responses. Across four tasks involving these inputs, we find that current baseline VLAs score under 3% accuracy in all but one case. GUI agents also struggle with this task, completing at most 3.9% of fields in the hardest Doc Transfer task. Among the steps involves in form-filling, we find that VLAs primarily struggle with text placement. GUI agents struggled with text placement, mulit-step actions, and completion within the allotted time frame.
To address the localization bottleneck, we introduce a modular architecture that separates semantic understanding from spatial grounding. Specifically, we equip any VLA with the ability to name the field it intends to complete, e.g., “Date of Birth”, and delegate the task of locating the corresponding input area to an auxiliary VLM FieldFinder tool. FieldFinder predicts the bounding box of the target field’s input region (e.g., an empty line, cell, check box, or empty space next to the target text). VLAs, when equipped with FieldFinder, improve accuracy by as much as 54 percentage points.
Our contributions are as follows:
@@
A benchmark for evaluating agents on realistic form completion scenarios, showing that current VLAs struggle to accurately identify field placements.
@@
An open-vocabulary field detection model, showing that it helps VLAs overcome spatial reasoning limitations.
We intend to release both publicly on GitHub.
Figure 1: The FormGym task. Agents are provided with a user profile in natural language and, optionally, a source form. The agent must use an editor API to complete the target form.
@@ 2 Related Work
Several benchmarks exist for evaluating document layout understanding Zhong et al. (2019); Pfitzmann et al. (2022); Li et al. (2020, 2019); Harley et al. (2015); Li et al. (2019). Numerous vision-language Xu et al. (2020); Li et al. (2021); Bao et al. (2020); Appalaraju et al. (2021); Lee et al. (2022) have been proposed for these types of tasks. Unlike traditional QA-style benchmarks, VLA evaluations generally measure a path-independent end-state, such as in Zhou et al. (2023), Zheng et al. (2022), Liu et al. (2023), Yao et al. (2024), and He et al. (2024), which often include elements of form-filling. Existing software, such as Mac OS Preview and Amazon Textract, can localize text fields in PDFs. However, they sometimes fail to identify non-underlined fields, including table cells or those indicated merely by a colon (e.g., "Name:  "). In contrast, our work builds on these domains to explore end-to-end, real-world form completion.
3 FormGym: Realistic Form-Filling for Agents
We aim to evaluate whether VLAs can produce completely filled forms when given access to user data and image editing tools. FormGym includes a diverse set of forms, user profiles, and agent actions representing a range of realistic challenges.
3.1 Documents
Our task consists of four document tasks. The Auto Loan Task - Text task consists of four densely annotated American vehicle loan application forms containing a total of 357 input fields. To enable evaluation on multiple user profiles (see below), we annotate each field with the type of user information (e.g., full name) it should contain rather than a specific answer (e.g., John Doe). For each form, we provide four user profiles. User profiles contain atomic facts, such as first name and postal code. As a result, many fields, such as address or middle initial, do not map directly to user profile information and instead must be derived from one or more user profile facts. In the case of the Auto Loans - Doc Transfer task, we provide the facts in the form of another Auto Loans source document, densely completed with user information. Information not available in the source document is provided in natural language.
The Database Task consists of 49 fields on two commercial banking forms. We provide the content of 39 of these fields in a SQL database that agents must query. Several of these fields are not provided in the SQL database so must be calculated arithmetically from values in other fields according to instructions on the form.
Finally, we contribute the FUNSD Task for evaluating diverse formats and multilingual reasoning, derived from Jaume et al. (2019)’s document relation dataset. The FUNSD Task consists of 50 examples from the FUNSD test set with exactly one target answer field masked in each document.
3.2 Actions
To edit forms, we provide agents with the following actions:
  • PlaceText(x, y, value) Place the text value centered at the coordinates $(x,y)$.
  • DeleteText(x, y) Delete all input text whose bounding boxes contain the coordinate $(x,y)$.
  • SignOrInitial(x, y, value) Place the value at coordinate $(x,y)$ in the form of a signature or initials.
  • QuerySql(query) Query the SQL database in the Database Task using query.
  • Terminate() End the current session.
3.3 Flows
We evaluate agents under two workflows:
@@ 3.3 Flows
Iterative - The agent may take multiple sets of actions over the course of up to 10 rounds, allowing it to correct mistakes. We report additional details in Appendix A.2.
3.4 Evaluation

### v3

@@ Abstract
End-to-end form filling refers to automatically populating fields in a document-style form with the appropriate information derived from external data. Although prevalent and useful, no formal benchmark exists for evaluating systems’ form completion accuracy. Existing datasets focus on parsing, extraction and web form interaction, rather than end-to-end completion of document-style forms. We propose FormGym, a benchmark formulation of the end-to-end form filling task that evaluates form completion and accuracy. We construct FormGym by repurposing three existing datasets and add one new dataset to achieve more challenging, diverse, and realistic test cases. Our studies show baseline vision language agents (VLAs) perform poorly on FormGym in every scenario, primarily due to poor field localization. GUI agents perform better but suffer from high latency and costs. Therefore we also introduce FieldFinder, a field localization tool that enables zero-shot VLAs to find and accurately place text in input fields. We find that VLAs augmented with FieldFinder achieve better performance compared to baselines in all models.
@@
Matthew Toles1, Isaac Song2,3, Rattandeep Singh1, Zhou Yu1,3, 1Columbia University, 2Georgia Institute of Technology, 3Arklex.ai, Correspondence: mt3639@columbia.edu
@@ 1 Introduction
Filling out paperwork is a pervasive and tedious task. The US government estimates that federal agencies collectively generate nearly 10 billion hours of mandatory paperwork each year USOMB (2016). Ad-hoc tools and advances among digital agents in related fields have shown potential to substantially reduce this burden Ghosh et al. (2024). However, we are unaware of any formal benchmark to evaluate automatic form filling. This results in uncertainty and risk for users, especially on forms carrying legal or financial weight.
Existing literature addresses components of end-to-end (E2E) form filling, including document understanding, question answering, tool use, and image manipulation. Work on GUI agent systems and benchmarks has primarily targeted browser and other desktop UIs. However, we are concerned with printable, PDF-style images of target documents. Lacking the same appearance and affordances as GUIs, these documents will challenge existing strategies, especially systems reliant on underlying structures such as the DOM or metadata.
In form filling tasks, each target document contains one or more empty fields, each associated with a field name (e.g. Occupation) in the document indicating the desired field value (e.g. Researcher) through layout semantics (e.g. proximity) (Figure 1). Field values are derived from an external source persona, such as free text or an image of a related document. We formalize the E2E form filling task mapping unfilled forms and source persona to filled forms as follows: for each field in the unfilled form, place an appropriate string inside its bounding box.
Figure 1: In the FormGym task, agents are provided with an unfilled source form and a user persona containing answers to fields in the form. The agent must use an editor API to produce the completed source form. Diverse layout semantics indicate suggested fields, such as underlines, colons, check boxes, and table cells.
We introduce FormGym, a novel benchmark dataset and evaluation framework for end-to-end form filling. To construct FormGym, we leverage three existing datasets originally constructed for document understanding. From these, we constructing ground truth field (name, bounding box, value) tuples for empty forms by redacting values. However, because of their original annotation schemes, it is difficult to construct realistic source personas without the field values appearing verbatim. We therefore extend our benchmark by manually annotating the Auto Loans dataset with a more sophisticated schema. This allows FormGym to address longer forms, incorporate secondary source documents into source personas, and include fields that require composing values from multiple facts rather than copying text directly.
Noting near-zero accuracy by baseline vision language agents (VLAs), we identify field localization as the main bottleneck in end-to-end form filling, even in frontier models. Rather than train purpose-built agents, we would prefer to provide flexible infrastructure adaptable to any new LLM. Hence, we develop FieldFinder, an open vocabulary model trained to output the bounding box of the field associated with an input field name. As a lightweight tool, FieldFinder allows VLAs to specify the target field by name rather than cartesian coordinates. We demonstrate that FieldFinder improves VLA form filling accuracy across all scenarios and models with minimal latency and memory overhead (Figure 1). Our contributions are twofold:
@@
  • A benchmark for evaluating agents on end-to-end form completion, showing that current VLAs and Claude Computer Use struggle to accurately identify field placements.
@@
  • An open-vocabulary field localization model, showing that it helps VLAs overcome spatial reasoning limitations.
We will release our dataset and code on GitHub after publication.
@@ 2 Related Work
Several benchmarks exist for evaluating document layout Zhong et al. (2019); Pfitzmann et al. (2022); Li et al. (2020, 2019); Harley et al. (2015); Li et al. (2019). Numerous frameworks Xu et al. (2020); Li et al. (2021); Bao et al. (2020); Appalaraju et al. (2021); Lee et al. (2022); Lu et al. (2024) have been proposed for navigating these types of tasks (Table 3). Despite deep exploration of document understanding, prior work has not addressed document elements that are suggested only by negative space in the document (e.g., Occupation in Figure 1). Some challenging examples include non-underlined fields, table cells or those indicated merely by a colon (e.g., "Name:  "). Existing commercial software, such as Mac OS Preview and Amazon Textract, can localize some but not all suggested elements.
Unlike traditional question answer-style (QA) benchmarks, vision language and graphical user interface (GUI) agent evaluations generally measure a path-independent end-state, such as in Zhou et al. (2023), Zheng et al. (2022), Liu et al. (2023), Yao et al. (2024), and He et al. (2024), which often include elements of form-filling. GUI agents including Qin et al. (2025b); Gou et al. (2025) generally focus training on GUIs rather than flat document images and interact with documents through browser UIs rather than editor APIs. In contrast, our work explores end-to-end, real-world completion of PDF-style image domain forms.
Computer use agents including Claude Computer Use Anthropic (2024) and Openai Operator OpenAI (2025) show potential to address these types of tasks. Open source frameworks have also been proposed, including Shen et al. (2024a); Qin et al. (2025a); Liu et al. (2024); Hong et al. (2024); Wei et al. (2025); Qi et al. (2025); Ma et al. (2023); Putta et al. (2024); Shen et al. (2024b). However, many of these solutions exploit structured elements such as the DOM, or are limited in their domain or interaction methods.
3 FormGym Benchmark
The FormGym dataset is composed of documents from three existing datasets plus one novel dataset. While existing datasets represent a large quantity of diverse, multilingual documents, our Auto Loans dataset introduces some of the most challenging scenarios in the form filling domain.
3.1 Conversion from Existing Datasets
We draw examples from the FUNSD Jaume et al. (2019), XFUND Xu et al. (2022), and Form-NLU Ding et al. (2023) document relation datasets, representing scanned English, scanned multilingual, and digital-born English financial documents, respectively (Figure 2). We choose these datasets because they contain a large quantity of relationship annotations between document field names and values, as well their bounding boxes (Table 1). FUNSD is composed scanned documents from the Truth Tobacco Industry Document archive University of California, San Francisco . XFUND is composed of scanned documents from Common Crawl in seven languages Common Crawl (2023). Form-NLU is composed of Australian financial filings. For each key-value relation, we create one (document image, field name, field value, value bounding box) example. We extract names, values, and bounding boxes from document annotations. On these documents, the correctness of predicted field values is determined by exact match. Because these datasets include only completed documents, we then create empty fillable documents by deleting value text from filled fields with horizontal inward content-aware fill.111github.com/light-and-ray/resynthesizer-python-lib
Forms Fields Fields/Form Sources Lang. Domain Quality Comp.
Train Test Train Test Train Test
Auto Loans - 10 - 886 88.6 4 1 Auto Loans Digital
FUNSD 155 39 2,246 577 14.5 14.8 1 1 US Government Scanned
XFUND 1,112 100 19,559 1,950 17.6 19.5 1 7 Common Crawl Scanned
Form-NLU 442 66 3,661 476 8.3 7.2 1 1 AU Stock Exchange Digital
Total 1,709 215 25,466 3,889 14.9 18.1
Table 1: FormGym Dataset statistics. Sources indicates the number of source personas for field value and prompt generation. Lang. indicates the number of languages. Comp. indicates whether field values must be generated by composing from multiple facts rather than merely copying a span.
3.2 Creating A New Auto Loans Dataset
While the above documents present a diversity of forms, the automatic nature of example creation does not explore the full complexity of real-world form filling. Specifically, field names and values may be phrased differently from the source persona (e.g. "previous address" vs. "last place of residence") and may require generation based on multiple facts, such as a street, city, and postal code. In addition, forms in existing datasets contain relatively few fields per form, substantially reducing the challenge of field localization. Finally, users may wish to have forms completed based on the content of other document images to avoid transcribing relevant information. In this case, we will require multiple forms containing partially overlapping field values. Therefore, we construct the additional Auto Loans dataset.
(a) Auto Loans
(b) FUNSD
(c) XFUND
(d) Form-NLU
Figure 2: Example forms and field bounding boxes in the FormGym dataset.
We searched the web for American auto loan application PDF files, choosing the single page with the most fields. One author then manually annotated the bounding box of each empty field. Another author reviewed these annotations and together they discussed and corrected any disagreements. We present additional annotation details in (Appendix A). We defined multiple source personas ensuring that all information necessary for every form be present, such as the applicant’s first, middle, and last name. This increases the diversity of source personas and enables better coverage of certain form fields, including check boxes, which are are sparsely filled. Observing that most enumerated options in these fields have four or fewer choices, we construct four personas, each spanning all documents. Finally, for each field, we define a function that determines prediction correctness based on one or more source persona elements. For example, the Full Name field will be marked as correct iff input == source_data.firstname + " " + source_data.lastname.
Defining source personas independently from form fields allows several forms of flexibility not available in XFUND, FUNSD, and Form-NLU. Specifically, it allows us to accept a broader range of inputs, for example, different phone number formats. It also permits automatically propagating other documents to serve as ground truth source persona inputs for the Auto Loans (Image) task (Section 3.2). By manually annotating empty forms instead of automatically redacting completed ones, we avoid artifacts from content-aware fill and biases from original text placement. This yields a subset with maximal evaluation precision.
3.3 FormGym Dataset Statistics
In total, FormGym contains 25,466 train and 3,889 test field examples spread across 1,709 train and 215 test forms (Table 1). The dataset contains between 7.2 and 88.6 fields per form with an average of 18.1. 46% of test documents are in one of seven non-English languages. 65% of documents are scanned versus digital-born.
4 FieldFinder Field Localizer
Figure 3: In the baseline case, the LLM receives an unfilled source form and persona information in its context and attempts to complete the form through a text placement API based on (x, y) coordinates. In the FieldFinder (ours) case, the text placement API is replaced with the FieldFinder tool. Instead of using coordinates as input, the FieldFinder tool takes the name of a field as input, then uses an open vocabulary object detection model to detect the corresponding field boudning box. In the GUI Agent case, the GUI agent uses an in-browser image editing tool (designed for humans) to place text on the PDf.
We find that baseline VLAs perform poorly, primarily due to poor field localization. We therefore design a field localization tool called FieldFinder that augments any VLA by taking on the localization task. FieldFinder is trained to take a form image and text description of the name of the target field as input and predicts the bounding box around the valid input space (Figure 3). Functionally, FieldFinder is an open-vocabulary object detection model, where the target objects are suggested text input fields associated with nearby text that matches the input string.
4.1 FieldFinder Dataset
We draw examples from the FUNSD, XFUND, and Form-NLU train datasets (Figure 2). In some cases, field names can only be described precisely by including additional hierarchical information, such as section headers. For example, in several Auto Loans forms, personal information for the applicant and the coapplicant are only indicated based on a distant section header. Similarly, specific table cells must be described by row, column, and table headers. We therefore prepend descriptions of all hierarchically superior elements to the input text (i.e., Section 1 | Members Table | Names Column | Row 1 $\rightarrow$ John Doe ). Examples consist of (field name, field bounding box) pairs such that FieldFinder can learn to localize the bounding box of a field’s value given its name.
4.2 FieldFinder Training
We fine-tune a Florence 2 Large Xiao et al. (2024) vision foundation model to predict the field bounding box given the target field name string and form image. We choose Florence 2 because its pretraining contains both open-vocabulary object detection and tasks requiring OCR, minimizing the distribution shift between pretraining and fine-tuning. Florence 2 Large has only 0.77B parameters, contributing minimal latency and memory overhead when augmenting with much larger VLAs. We train the FieldFinder for 6 epochs, batch size 8, learning rate 1e-6 with cosine profile on 1x NVIDIA A100 GPU for 30 hours. We selected these parameters using grid search across 2-4 options per parameter.
4.3 FieldFinder Results
We choose accuracy as our primary metric, defined as whether the center of the predicted bounding box falls within the ground truth bounding box. FieldFinder achives 54.3% accuracy on average across the the test sets of all FUNSD, XFUND, and Form-NLU, compared to near-zero accuracy by baseline VLAs. However, we note a wide disparity in performance depending on the dataset (Figure 4). We see the strongest performance in Form-NLU at 80.5% accuracy, which we attribute to its few fields per form and lack of scan artifacts. Accuracy decreases on FUNSD to 57.4%, which contain low quality scans. XFUND presents an even greater challenge due to its multilingual scans, with FieldFinder reaching 24.9% accuracy. Finally, we find weak performance on Auto Loans at only 6.9% accuracy, which did not appear in the training data and contains by far the most fields per form, at 88.6 fields per form. We conclude that FieldFinder can perform well on high quality English documents with few fields per form, but struggles in other contexts and out-of-distribution data.
Figure 4: FieldFinder accuracy vs. fields per form (log scale). Trend line shown for English datasets only. The trend suggests that high numbers of fields per form and multi-lingual forms are the greatest challenges for FieldFinder.
Although distribution shift likely affects quality on Auto Loans, we draw attention to a confounding factor; Auto Loans documents are so densely covered with fillable fields (88.6 per page) that there is a high probability for an errant placement to land in an unrelated field, resulting in two errors—an empty filed and a doubly filled field. When accuracy is plotted against fields per form on a log scale, English datasets Form-NLU, FUNSD, and Auto Loans shows a clear negative linear trend (Figure 4). This further indicates that FieldFinder accuracy highly dependent on the number of fields per page.
5 End-to-End Form Filling Pipelines
Noting the strong performance of VLAs on related visual reasoning and QA tasks, we instantiate an E2E form filling VLA pipeline for their evaluation. While many methods have been proposed for related tasks, only VLAs possess the necessary capabilities to fully perform end-to-end form filling (Appendix C). We construct parallel pipelines for VLAs augmented with our FieldFinder tool and Claude Comptuer Use GUI agent.
5.1 VLA Pipeline
We choose leading open- and closed-source VLAs as our baselines because they represent the only end-to-end solutions that the necessary criteria in table 3. We select Llava 7B, Molmo 7B, and Aria 25B as our open-source; and Claude 4 and GPT-5 as our frontier VLAs. We prompt VLAs with the source image, source persona and form editing API documentation. We also evaluate Claude 4 + Set of Marks Yang et al. (2023), overlaying a grid with coordinate references at vertices.
To leverage VLAs existing tool use capabilities, we construct a minimal image editing API by which they can edit the form. The image editing API permits placing text or signatures/initials on the image by specifying the string and (x, y) coordinates. It also allows for deleting all input text intersecting an (x, y) coordinate and terminating a session early.
5.2 FieldFinder Pipeline
We evaluate VLAs augmented with FieldFinder in a pipeline consistent with the previously described VLA pipeline. However, we replace the text and signature placement APIs with a FieldFinder API. Rather than taking Cartesian coordinates as input along with the field value string, the FieldFinder API takes a field name and value. It then places the value at the location detected by FieldFinder based on the field name.
5.3 GUI Agent Pipeline
We instantiate the Claude Computer use Use GUI agent with the free in-browser photo editing application Photopea222photopea.com, whose interface is nearly identical to Adobe Photoshop. We prompt Claude Computer Use with the natural language source persona and instructions to complete the form. Prompts include detailed instructions on how to use the Photopea interface, without which GUI agents fail completely. We provide additional implementation details in Appendix E.1. For accessibility and cost reasons, we limit operators to five minutes per page and downsample test sets to 10 documents per source.
AL (Text) AL (Image) FUNSD XFUND Form-NLU Average Cost
OS IT OS IT OS OS OS
Aria 25B 0.0 0.0 0.0 0.0 1.0 1.0 0.0 0.4 -
Claude 4 0.0 0.0 0.3 0.7 21.0 1.0 0.0 4.5 0.76
GPT-5 1.0 1.5 0.5 0.7 2.0 2.0 0.0 1.2 0.30
Llava 7B 0.0 0.0 0.0 0.0 1.0 1.0 0.0 0.4 -
Molmo 7B 0.0 0.0 - - 1.0 1.0 0.0 0.5 -
Claude 4 + SoM 0.0 1.0 1.0 1.0 9.0 1.0 0.0 2.3 0.69
Aria 25B + FF (ours) 3.3 4.2 1.5 2.0 20.0 9.0 29.0 12.7 -
Claude 4 + FF (ours) 8.3 9.2 4.8 5.3 32.0 15.0 54.0 23.0 0.43
GPT-5 + FF (ours) 8.5 9.8 3.0 5.3 29.0 14.0 50.0 21.3 0.37
Llava 7B + FF (ours) 1.0 1.0 0.3 0.3 4.0 3.0 4.0 2.5 -
Molmo 7B + FF (ours) 0.5 1.3 - - 9.0 3.0 7.0 5.0 -
Claude Computer Use 2.7 - 1.4 - 46.2 26.7 32.4 21.1 54.00
Claude 4 (GT Coords) 74.0 90.0 49.2 56.2 83.0 82.0 77.0 75.3 -
Table 2: Percentage of fields containing a correct value. AL: Auto Loans dataset. IT: Iterative. OS: One-Shot. $F$: FieldFinder. SoM: Set of Marks. GT: VLA was supplied with ground truth field coordinates in the prompt. Macro average is taken across each dataset. Cost indicates API fees in USD per thousand fields. We do not include Molmo in the Auto Loans (Image) task because it is not pretrained with multiple images.
6 Experiments and Results
6.1 Pipeline Settings
Because Auto Loans documents contain numerous fields per page, we evaluate VLAs under two settings:
@@ 6.1 Pipeline Settings
Iterative - The agent may take multiple sets of actions over the course of up to 5 rounds, allowing it to correct mistakes. Agents receive feedback on the success or failure of each action.
The Iterative setting allows agents to adapt to tooling and to realize and correct mistakes using the deletion API endpoint. The GUI agent pipeline is inherently iterative. We present source personas in two forms:
Text - The source persona is presented as plain text.
Image - The source persona is presented as an image of another auto loan document. Persona data not appearing in the image is presented in plain text.
We evaluate VLAs in our E2E form filling pipeline with and without FieldFinder on the FormGym dataset. We find FieldFinder improves accuracy in all cases.
6.2 Metrics